OpenStudy (anonymous):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (imnotgoodatmath):

a kilobyte is 2^10 bytes you're converting 1 megabyte to kilobyte. short answer: divide 1 megabyte by 2^10. 2^20 / 2^10 = 2^10 kilobyte long answer: converting units that you learned in science class (1 megabyte / 1) * (2^20 bytes / 1 megabyte) * (1 kilobyte / 2^10 bytes) megabyte and bytes cancels out and you're left with 2^20 kilobyte / 2^10 = 2^10 kilobyte